# Policy Optimization in a Noisy Neighborhood: On Return Landscapes in Continuous Control

### v1

@@ 1 Introduction
Noisy neighborhoods in the return landscape. We call the return landscape the mapping from $\bm{\theta}$ to $R(\bm{\theta}$), our main object of study. We show that the return often varies substantially within the vicinity of any given $\bm{\theta}$, forming what we call a noisy neighborhood of $\bm{\theta}$.  Based on this observation, we demonstrate the usefulness of studying the landscape through the distribution of returns obtained from small perturbations of $\bm{\theta}$. In the important case where these perturbations result from a single gradient step, we call the resulting object the post-update return distribution.
@@
We believe the phenomenon we study is central to deep reinforcement learning in continuous control. Beyond its effect on learning dynamics (for example, through increased variance and implicit exploration (Schaul2022ThePO, )), it is also a potential driver of instability in sim2real settings, even in the face of seemingly small environmental changes. Additionally, it suggests that one should not simply deploy the policy obtained at the end of a training run, and that further post-training tuning may be beneficial.
@@ 2 Background
We are interested in understanding how small changes to the policy parameter affect the associated return. To this end it is sufficient to study the return from the reference state $s_{I}$ (in Appendix A.4 we show that similar effects occur across the state space). The return landscape is our main object of study.
@@ 3.2 Analyzing Failures
One characteristic feature of the post-update distributions studied above is the existence of a significant lower tail for many policies visited by the three deep RL algorithms TD3, SAC and PPO. This is visible in their skewness, but especially in their left-tail probability, which demonstrates that many policies produce returns which are unexpectedly poor after up to roughly 10% of updates. We now take a closer look at the specific mechanism by which small changes in an agent’s actions result in a wide range of returns in continuous control.
@@ 4.1 Connectivity in the Return Landscape
For our analysis, we use 200 policies generated by different runs of TD3. From these we select pairs of policies with different post-update return distributions, as measured by their standard deviation or left-tail probability, but similar mean. Consider two sets of policy parameters $\bm{\theta}_{1}$ and $\bm{\theta}_{2}$, for which the post-update return distribution implied by $\bm{\theta}_{1}$ has higher LTP than that implied by $\bm{\theta}_{2}$. We linearly interpolate these two to form a family of parameters $\bm{\theta}=\alpha\bm{\theta}_{1}+(1-\alpha)\bm{\theta}_{2},\alpha\in[0,1]$. For each such $\bm{\theta}$, we then record the return $R(\bm{\theta})$ obtained by a single simulation with the corresponding policy.
In Figure 5, we show the result of this interpolation for six pairs of policies in the hopper and walker2d environments, in two distinct cases. In the first case, the two policies have been produced by the same run of TD3 (i.e., starting from the same initialization and history of batches); in the second case, the two policies have been generated by independent repetitions of the algorithm. The plot shows interesting information about the global structure of the return landscape: the interpolation process traverses different parts of the landscape, highlighting a transition between a noisy part of the landscape to an inherently smoother one. Interestingly, the interpolations between policies from the same run and those from different runs exhibit very different qualities. When interpolating between policies of different runs, the process traverses entire regions of the landscape of poor return, until the point in which it gets to the neighborhood of the second policy. By contrast, when interpolating between policies from the same run, the transition from a noisy to a smooth landscape happens without encountering any valley of low return – even when these policies are separated by hundreds of thousands of gradient steps in training. This is particularly surprising given that $\bm{\theta}$ is a high-dimensional vector containing all of the weights of the neural network, and there is no a priori reason to believe that interpolated parameters should result in policies that are at all sensible.
@@
To further quantify the phenomenon, we want to measure the proportion of return collapses encountered when interpolating between policies. We use the following experimental design. We sample for each environment a set of 500 pairs of policies from the same runs and a set of 500 pairs of policies from different runs. Then, we linearly interpolate between policies in the pairs, producing 100 intermediate policies, and randomly perturb them using Gaussian noise with standard deviation $3\times 10^{-4}$ to obtain an estimate of the mean of their (random) post-update return distribution. Then, for each pair of policies, we measure how frequently the mean return collapses in between the two extremes, by counting how many times it becomes less than 10% of the minimum return of the two original policies. We then average this Below-Threshold Proportion across pairs, and across environments using rliable (Agarwal2021DeepRL, ). Figure 6 shows that there is on average almost no drop in return when interpolating among policies from the same run. We additionally report similar results on four ALE games in Appendix A.6.
@@ 4.2 Stabilizing Policies by Navigating the Landscape
In Figure 7, we compare the improvements in LTP that these two variants induce when applied to the same policy, aggregated across Brax tasks, using at least 10 policies per environment, after only forty gradient steps. We additionally present scatter plots demonstrating the effect of applying Algorithm 1 to individual policies in Appendix A.10. Our results demonstrate that our Algorithm 1, by leveraging rejection, can be an effective tool for reducing the LTP of an existing policy.

### v3

@@ 1 Introduction
Noisy neighborhoods in the return landscape. We call the return landscape the mapping from $\bm{\theta}$ to $R(\theta$), our main object of study. We show that the return often varies substantially within the vicinity of any given $\bm{\theta}$, forming what we call a noisy neighborhood of $\bm{\theta}$. Based on this observation, we demonstrate the usefulness of studying the landscape through the distribution of returns obtained from small perturbations of $\bm{\theta}$. In the important case where these perturbations result from a single gradient step, we call the resulting object the post-update return distribution.
@@
We believe the phenomenon we study is central to deep reinforcement learning in continuous control, a finding which parallels a sensitive dependence on initial conditions observed previously in the Cartpole domain (parmas2018pipps, ). Beyond its effect on learning dynamics (for example, through increased variance and implicit exploration (Schaul2022ThePO, )), it is also a potential driver of instability in sim2real settings, even under small environmental changes. Additionally, it suggests that one should not simply deploy the policy obtained at the end of a training run, and that further post-training tuning may be beneficial.
@@ 2 Background
We are interested in understanding how small changes to the policy parameter affect the return. To this end it is sufficient to study the return from the reference state $s_{I}$ (in Appendix A.4 we show that similar effects occur across the state space). The return landscape is our main object of study.
@@ 3.2 Analyzing Failures
One characteristic feature of the post-update distributions studied above is the existence of a significant lower tail for many policies visited by the three deep RL algorithms TD3, SAC and PPO. This is visible in their skewness, but especially in their left-tail probability, which demonstrates that many policies produce returns which are unexpectedly poor after up to roughly 10% of updates. We now take a closer look at the specific mechanism by which small changes in an agent’s actions results in a wide range of returns in continuous control.
@@ 4.1 Connectivity in the Return Landscape
For our analysis, we use 200 policies generated by different runs of TD3. From these we select pairs of policies with different post-update return distributions, as measured by their standard deviation or left-tail probability, but similar mean. Consider two sets of policy parameters $\bm{\theta}_{1}$ and $\bm{\theta}_{2}$, for which the post-update return distribution implied by $\bm{\theta}_{1}$ has higher LTP than the implied by $\bm{\theta}_{2}$. We linearly interpolate these two to form a family of parameters $\bm{\theta}=\alpha\bm{\theta}_{1}+(1-\alpha)\bm{\theta}_{2},\alpha\in[0,1]$. For each such $\bm{\theta}$, we then record the return $R(\bm{\theta})$ obtained by a single simulation with the corresponding policy.
In Figure 5, we show the result of this interpolation for six pairs of policies in the hopper and walker2d environments, in two distinct cases. In the first case, the two policies have been produced by the same run of TD3 (i.e., starting from the same initialization and history of batches); in the second case, the two policies have been generated by independent repetitions of the algorithm. The plot shows interesting information about the global structure of the return landscape: the interpolation process traverses different parts of the landscape, highlighting a transition between a noisy part of the landscape to an inherently smoother one. Interestingly, the interpolations between policies from the same run and from different runs exhibit very different qualities. When interpolating between policies of different runs, the process traverses entire regions of the landscape of poor return, until the point in which it gets to the neighborhood of the second policy. By contrast, when interpolating between policies from the same run, the transition from a noisy to a smooth landscape happens without encountering any valley of low return – even when these policies are separated by hundreds of thousands of gradient steps in training. This is particularly surprising given that $\bm{\theta}$ is a high-dimensional vector containing all of the weights of the neural network, and there is no a priori reason to believe that interpolated parameters should result in policies that are at all sensible.
@@
To further quantify the phenomenon, we want to measure the proportion of return collapses encountered when interpolating between policies. We use the following experimental design. We sample for each environment a set of 500 pairs of policies from the same runs and a set of 500 pairs of policies from different runs. Then, we linearly interpolate between policies in the pairs, producing 100 intermediate policies, and randomly perturb them using Gaussian noise with standard deviation $3\times 10^{-4}$ to obtain an estimate of the mean of their (random) post-update return distribution. Then, for each pair of policies, we measure how frequently the return collapses in between the two extremes, by counting how many times it becomes less than 10% of the minimum return of the two original policies. We then average this Below-Threshold Proportion across pairs, and across environments using rliable (Agarwal2021DeepRL, ). Figure 6 shows that there is on average almost no drop in return when interpolating among policies from the same run. We additionally report similar results on four ALE games in Appendix A.6.
@@ 4.2 Stabilizing Policies by Navigating the Landscape
In Figure 7, we show the improvement in LTP that this algorithm induces when applied to the same policy, aggregated across Brax tasks, using at least 10 policies per environment, after only forty gradient steps. We additionally present scatter plots demonstrating the effect of applying Algorithm 1 to individual policies in Appendix A.10. Our results demonstrate that this rejection procedure can be an effective tool for reducing the LTP of an existing policy.